OpenStudy (anonymous):

x^2+y^+3x+4y+4=0 How do you get the standard form of the circle?

OpenStudy (anonymous):

you must complete the square for both x and y x^2 + 3x = (x + 3/2)^2 - 9/4 [if you expand it you'll see it's true] y^2 + 4y = (y + 2)^2 - 4 whole thing is now (x + 3/2)^2 - 9/4 + (y + 2)^2 - 4 + 4 = 0 (x + 3/2)^2 + (y + 2)^2 = 9/4 if you're unclear about completing the square: (x + a)^2 = x^2 + 2a + a^2 so x^2 + 2a = (x + a)^2 - a^2